# World-to-Words: Grounded Open Vocabulary Acquisition through Fast Mapping in Vision-Language Models

### v1

@@ 1 Introduction
Language is learned through sensorimotor experience in the physical world (Bisk et al., 2020). The ability to connect language units to their referents in the physical world, referred to as grounding, plays an important role in learning and understanding grounded meanings of words (Harnad, 1990). As shown in Figure 1, a human reader would easily ground noun phrases to the corresponding entities captured in the image. Even when the term “incinerator” is new to human learners, they can still locate the object of interest through the language and visual context, and acquire its meaning. In fact, this ability to bootstrap new word learning with only minimal information, known as fast mapping, is demonstrated abundantly in cognitive literature on human language acquisition (Carey and Bartlett, 1978; Carey, 1978; Golinkoff et al., 2000; Smith and Yu, 2008).
@@ 5 Related Work
Vision-Language Mapping
Mapping plays a central role in classic lexicon acquisition problem (Gleitman and Landau, 1994; Clark, 1995). Primarily, researchers focused on grounding words to their meaning symbols, building learning mechanisms using specific mental biases to simulate children’s word acquisition, and giving computational accounts for psycholinguistic phenomena (Siskind, 1996; Regier, 2005; Goodman et al., 2007; Fazly et al., 2010). Early efforts along this line incorporate visual grounding either by learning a statistical or neural mapping from object categories (Roy and Pentland, 2002; Yu, 2005; Xu and Tenenbaum, 2007; Yu and Ballard, 2007; Yu and Siskind, 2013) and more complicated visual features (Qu and Chai, 2010; Mao et al., 2019, 2021; Pratt et al., 2020) to linguistic labels. These studies are usually in a closed world with limited vocabulary (Krahmer and van Deemter, 2019), and words are usually isolated from the natural context of use. More recently, there exist multi-modal understanding tasks, e.g., object retrieval (Guadarrama et al., 2014; Hu et al., 2016), referring expression comprehension (Liu et al., 2014; Yu et al., 2016; Mao et al., 2016; Wu et al., 2020), and phrase grounding (Plummer et al., 2015) to map text to corresponding objects. Our setup is closely related to this line as we position grounding as an explicit word-referent mapping problem. The difference is that, our work goes beyond grounding to study open-vocabulary acquisition through fast mapping, a more complicated but realistic challenge faced by AI agents.
@@
To evaluate vision-language systems, many downstream tasks have been formulated. Some related formulations are summarized in Table A.1 in Appendix. While demonstrating some vision-language capabilities, these down-stream tasks provide limited insights into whether these models truly capture the grounded meaning of words with respect to the external environment. Our task design specifically targets the machine’s ability to predict words and ground words to perception. More akin to our formulation is the vision-based language modeling task (Jin et al., 2020) in a continual learning setting. Our work differs mainly in two aspects. First, the proposed task only predicts masked tokens based on the visual context, which leaves the referential uncertainty (i.e., grounding) unattended (e.g., in Figure 2, correct prediction of the word “boat” does not guarantee correct grounding). Also, Jin et al. (2020) focuses on compositionality, while we seek to address few-shot grounded word learning when unseen words are encountered after pre-training.

### v2

@@ 1 Introduction
Language is learned through sensorimotor experience in the physical world (Bisk et al., 2020). The ability to connect language units to their referents in the physical world, i.e. (referential) grounding, plays an important role in learning and understanding grounded meanings of words (Harnad, 1990). As shown in Figure 1, a human reader would easily ground noun phrases to the corresponding entities captured in the image. Even when the term “incinerator” is new to human learners, they can still locate the object of interest through the language and visual context, and acquire its meaning. In fact, this ability to bootstrap new word learning with only minimal information, known as fast mapping, is demonstrated abundantly in cognitive literature on human language acquisition (Carey and Bartlett, 1978; Carey, 1978; Golinkoff et al., 2000; Smith and Yu, 2008).
@@ 5 Related Work
Grounding Language to Vision
Referential grounding plays a central role in classic lexicon acquisition problem (Gleitman and Landau, 1994; Clark, 1995). Primarily, researchers focused on grounding words to their meaning symbols, building learning mechanisms using specific mental biases to simulate children’s word acquisition, and giving computational accounts for psycholinguistic phenomena (Siskind, 1996; Regier, 2005; Goodman et al., 2007; Fazly et al., 2010). Early research in this area focused on visual grounding through two main approaches: first, by developing statistical and neural mappings from object categories to lexical items (Roy and Pentland, 2002; Yu, 2005; Xu and Tenenbaum, 2007; Yu and Ballard, 2007; Yu and Siskind, 2013), and later by incorporating more sophisticated visual features to establish connections with lexicon, syntax, and complex semantic structures (Qu and Chai, 2010; Shi et al., 2019; Mao et al., 2019, 2021; Pratt et al., 2020). These studies are usually in a closed world with limited vocabulary (Krahmer and van Deemter, 2019), and words are usually isolated from the natural context of use. More recently, there exist multi-modal understanding tasks, e.g., object retrieval (Guadarrama et al., 2014; Hu et al., 2016), referring expression comprehension (Liu et al., 2014; Yu et al., 2016; Mao et al., 2016; Wu et al., 2020), and phrase grounding (Plummer et al., 2015) to map text to corresponding objects. Our setup relates to this line as we position referential grounding as an explicit word-referent mapping problem. We focus on efficient learning of grounded lexical semantics through fast mapping in modern vision-language models, a more scalable but realistic challenge faced by AI agents.
@@
To evaluate vision-language systems, many downstream tasks have been formulated. Some related formulations are summarized in Table A.1 in the Appendix. While demonstrating some vision-language capabilities, these downstream tasks provide limited insights into whether these models truly capture the grounded meaning of words with respect to the external environment. Our task design specifically targets the machine’s ability to predict words and ground words to perception. More akin to our formulation is the vision-based language modeling task (Jin et al., 2020) in a continual learning setting. Our work differs mainly in two aspects. First, the proposed task only predicts masked tokens based on the visual context, which leaves the referential uncertainty (i.e., grounding) unattended (e.g., in Figure 2, correct prediction of the word “boat” does not guarantee correct grounding). Also, Jin et al. (2020) focuses on compositionality, while we seek to address few-shot grounded word learning when unseen words are encountered after pre-training.
@@ A.1 Illustrated Comparison of Setting
A.2 Evaluation Protocols Explained
We present an adequate evaluation protocol for grounded word acquisition in the main paper. This section provides more in-depth explanation for the metrics and implementation details for reproducibility purposes.
Perplexity Metric Details
We follow prior practice in cloze tests (Salazar et al., 2020; Jin et al., 2020) to evaluate the perplexity of a word $w$. We use log pseudo-perplexity in masked language modeling, defined as
$\log\text{PPL}(w)=-\log P(w|x_{\text{img}},x_{\text{cap}})$
However, the majority of the language models employ sub-word tokenization methods to segment and encode text. In particular, one lexical word can be segmented into several tokens, and different tokenizers can lead to different tokens for the same input. We thus introduce a tokenizer-dependent measure for perplexity. For tokenizer $T$, we represent the $N$ tokens of word $w$ as $T(w)$ and
$\log\text{PPL}(w)=-\frac{1}{N}\sum_{t\in T(w)}\log P(t|x_{\text{img}},x_{\text{cap}})$
IoU Metric Details
we face the same challenge as Kamath et al. (2021) where multiple referents are possible for a masked word. In a similar manner, we adopt the Any-Protocol and All-Protocol to evaluate the grounded detection task. Assuming $n$ ground truth bounding boxes $B=\{b_{1},b_{2},\cdots,b_{n}\}$ and $m$ predicted bounding boxes $\widetilde{B}=\{\widetilde{b_{1}},\widetilde{b_{2}},\cdots,\widetilde{b_{m}}\}$. The intersection-over-union (IoU) under Any-Protocols is defined as the average IoU of the best matching predicted bounding box for each ground truth object:
$\text{IoU}_{\text{any}}=\frac{1}{n}\sum_{i\in\{1,2,\cdots,n\}}\max_{j\in\{1,2,\cdots,m\}}\text{IoU}(b_{i},\widetilde{b_{j}})$
The intersection-over-union (IoU) under All-Protocols is defined as the IoU between the joint bounding box of ground truth and predicted bounding boxes:
$\text{IoU}_{\text{all}}=\text{IoU}(\cup B,\cup\widetilde{B})$
A.3 Word List
  • 60 words are in the seen-set, each with 80 test cases: baby, ball, beach, bench, bike, black, blond, blue, boy, brown, building, car, child, dark, dog, dress, face, female, field, floor, food, girl, glasses, grass, gray, green, guitar, guy, hair, hand, hat, head, horse, jacket, jeans, lady, large, little, long, man, orange, pants, person, player, red, shirt, sidewalk, sign, small, snow, street, striped, table, top, wall, water, white, woman, yellow, young.
  • 31 words are in the unseen-set, each with 50 test cases222a few words (product, steep, telephone) has one less test case due to the availability of Flickr30K Entities Dataset.: aged, bamboo, barefoot, brush, button, cafe, cheese, circular, classroom, crosswalk, diverse, doctor, donkey, elephant, fluffy, foreign, gym, heart, newborn, pan, pizza, product, security, sink, star, steep, stove, student, teacher, telephone, warm.
Appendix B Computational Model Details
B.1 Pre-training Objectives
Masked Language Modeling (MLM).
The MLM head can be placed at multiple possible places, and our design is an exploration after preliminary experiments on smaller-scale training. We strictly follow the setup of RoBERTa to implement the MLM head with a two-layer MLP, based on the implementation of huggingface [^3] https://huggingface.co/docs/transformers/model_doc/roberta. Words in groundable phrases are masked with a probability of 0.4 and those in non-groundable regions are masked with a lower probability of 0.1. For a token selected to mask, we follow RoBERTa to assign a probability of 80% to replace with MASK, 10% with a random token, and 10% to do nothing.
Object Localization (OL).
We follow MDETR to decode object embeddings with a three-layer MLP to produce bounding boxes. Similar to most prior work, we apply a filter over boxes with confidence below 0.7. In our framework, this means that the object corresponds to the no-object label $\varnothing$ (Figure 4) with a probability over 0.3. We strictly follow DETR to perform bipartite matching between proposed boxes and ground truth boxes with a Hungarian loss. The predicted boxes are optimized towards ground truth by the generalized intersection-over-union (GIoU) loss and the L1 loss.
Grounding.
In positional alignment, the model learns to map each object representation to tokens in the sentence with a fixed length of 257, which could possibly be a MASK or an additional no-object label $\varnothing$ (Figure 4). The object and the token are considered a match given a mapping probability over 0.1. We use a fully-connected layer to predict the distribution over token positions with cross-entropy loss. In semantic alignment, the model learns to bring word embeddings closer to the object embeddings that they ground to, and push the unrelated pairs farther. We strictly follow the contrastive loss function defined in MDETR for every object and groundable token for this purpose.
B.2 Few-shot Learning Details.
Since no bounding box or word-object mappings annotation is available, we train OctoBERT with only masked language modeling (MLM) in few-sample new word learning. We reduce the batch size to 8 considering the fewer number of samples, and set the convergence criteria to a fixed number, i.e., 50 steps. All the rest of the experimental settings remain the same as pre-training.
Appendix C Experiment Reproducibility
C.1 OctoBERT Implementation Details
Our OctoBERT model mainly consists of one cross-modal transformer with inputs from uni-modal encoders from image and text domain. Specially, we select the ResNet-50 He et al. (2016) pretrained on ImageNet from TIMM [^4] https://github.com/rwightman/pytorch-image-models as the image encoder, and RoBERTa-base Liu et al. (2019) from huggingface [^5] https://huggingface.co/docs/transformers/model_doc/roberta as the text encoder. The cross-modal encoder and two decoders each consists of 4 transformer blocks with 8 attention heads, an input and output dimensionality of 512, and an inner-layer dimensionality of 2,048. Besides, 50 learnable object queries are included to query the cross-modal decoder to generate bounding box proposals.
C.2 Hyper-parameter Decisions
We include the major hyper-parameter tuning decisions for reproducibility purpose. For more details, please refer to the supplementary codes.
  • Learning Rate:
    • Image Encoder: frozen
    • Text Encoder: $1\times 10^{-5}$
    • Multi-modal Transformer: $1\times 10^{-4}$
  • Batch Size: 128
  • Pre-training Loss Coefficients:
    • MLM Loss: 32
    • Cross Entropy for Positional Alignment: 1
    • Contrastive Loss for Semantic Alignment: 1
    • L1 Localization Loss: 5
    • GIoU Localization Loss: 2
  • Few-shot Learning:
    • Batch size: 8
    • Other Hyper-parameters: Same as Pre-training
C.3 Computational Resources
Our OctoBERT models is pre-trained on 8 NVidia A40 GPUs. With mixed-precision pre-training and a batch size of 128, OctoBERT was trained for 150,000 steps where each step takes about 1.4 second.
C.4 Evaluation on GOVA
OctoBERT
For our proposed OctoBERT model, given a GOVA test, with its corresponding image and textual cloze pair passing into the model, the bounding box predictions are generated by keeping only the bounding box proposals that are mapped to at least one masked token within the cloze, while the masked token prediction results are directly decoded from its language modeling head.
VisualBERT
For the “Detect-and-Recognize” baseline model VisualBERT, we use phrase-grounding fine-tuned version of VisualBERT to perform object localization, and, as it lacks the language modeling head, another vanilla pre-trained VisualBERT to perform mask token prediction. Specifically, for the bounding box localization part, we treat it as a standard phrase grounding task and follow Li et al. (2019) to select the top-1 bounding box prediction in the last masked token as the output.
ViLT+MDETR
For the “Produce-and-Localize” baseline model ViLT + MDETR, in stage one, we feed the input image and text into ViLT, collecting its top-1 cloze token prediction result. Then, at stage two, the input image and ViLT-completed text are fed into MDETR, performing phrase-grounding to localize the object associated with the original cloze. Finally, the cloze token prediction result from ViLT together with the bounding box proposals from MDETR are used for GOVA evaluation.
Appendix D Addendum to Results
D.1 Ablation Study
We performed an ablation study on several OctoBERT model variants to pinpoint what makes our OctoBERT model effective. These included models without language encoder initialization (w/o Init), without grounding objective (w/o G), without any object-centric representation (w/o O), and a text-only setup without any vision input (w/o V). For consistency, we control the number of transformer layers and the number of parameters for each variation. Despite tweaking various hyperparameters, no significant improvements were observed. As a result, we retained the same hyperparameters as in the OctoBERT model.
  • w/o G: This refers to the model variant without grounding loss, as has already been described in Section 3.2;
  • w/o O: This variant excludes all object-centric representations, retaining only the masked language modeling (MLM) objective. With this model, the object decoder transformer is unnecessary, thus no grounding nor localization is performed. Instead, we consolidate all 12 transformer blocks into the multi-modal encoder and directly attach the MLM objective to it.
  • Text-Only: This text-only model operates without any vision input or supervision, reducing it to a unimodal language model (RoBERTa) with 12 additional transformer blocks.
Following the analysis of Chang and Bergen (2022) in unimodal language models, we present the KL-Divergence between the model predictions and the unigram distribution in Figure 9. An immediate observation is that all variants converge to the shallow unigram statistics at around $10^{2}$ steps of pre-training. This aligns with the findings of Chang and Bergen (2022) that unimodal language models would converge to unigram before acquiring more complicated contextual representations. We noticed that in both text-only and OctoBERTw/o ${}_{\textrm{w/o O}}$ cases where MLM is the only pre-training objective, the models tend to stay around the unigram word distribution even with $10^{4}$ steps of training. However, variants with an object-centric representation quickly departed from the unigram distribution. Comparatively, models with language model initialization move quickly away from the unigram distribution, and models with a grounded objective have a marginally faster deviation. These results confirm that vision-language models can benefit from unimodal pre-training on a large corpus, and that performing language modeling upon object representations is crucial. We note that we compare the KL-Divergence from unigram distribution only to understand the models’ behaviors, and the metric itself can not serve as an evaluation of a system’s performance in grounded open vocabulary acquisition.
Figure 9: KL-divergence between model’s token prediction and the unigram distribution of the training corpus.
D.2 Addendum to Results in Multi-Class Incremental Learning
We present additional results in Table 6.
[…]
0 1.79 2.33 11.58 11.89
8 3.15 3.63 3.09 3.32
16 3.36 3.76 2.64 2.85
24 3.05 3.46 2.07 2.67
32 3.07 3.62 2.01 2.54
Table 6: The log G-PPL (All-Protocol) of seen and unseen words in multi-class incremental learning, each unseen word with a sample size ranging from 8 to 32.
D.3 Learning New Words through One-Class Incremental Learning.
We further perform a more controlled study with a word-specific one-class incremental learning setting. The pre-trained model is tasked to acquire one single unseen word from a few-shot learning session with $|\mathcal{V}_{\text{unseen}}|=1$. The results of this section are obtained from the test immediately following the new session. We present the test result in Table 7. Again, we observe that with as few as 8 samples, OctoBERT can achieve a satisfyingly low grounded perplexity. In the majority of the cases, OctoBERT demonstrates the better ability to acquire unseen words over the groundless baseline.
# Samples 0 8 16 24 32 # Samples 0 8 16 24 32
crosswalk OctoBERT 10.82 8.48 7.43 7.70 5.95 donkey OctoBERT 8.70 0.84 0.81 0.67 0.79
OctoBERTw/o ${}_{\textrm{w/o G}}$ 10.91 10.88 7.53 7.15 7.5 OctoBERTw/o ${}_{\textrm{w/o G}}$ 9.69 1.97 1.99 2.35 2.01
cheese OctoBERT 12.16 2.62 3.00 1.27 1.04 barefoot OctoBERT 9.71 6.93 4.58 5.55 6.27
OctoBERTw/o ${}_{\textrm{w/o G}}$ 13.07 2.81 3.13 2.56 1.49 OctoBERTw/o ${}_{\textrm{w/o G}}$ 9.95 6.52 4.67 5.74 5.88
star OctoBERT 8.70 1.49 1.47 1.09 1.18 elephant OctoBERT 15.24 1.44 1.65 1.81 1.44
OctoBERTw/o ${}_{\textrm{w/o G}}$ 10.59 2.93 2.10 1.99 1.39 OctoBERTw/o ${}_{\textrm{w/o G}}$ 14.75 2.17 1.98 1.73 1.61
classroom OctoBERT 3.96 0.47 0.36 0.43 0.32 heart OctoBERT 9.34 2.97 1.90 1.76 1.76
OctoBERTw/o ${}_{\textrm{w/o G}}$ 5.10 0.95 0.88 1.05 0.95 OctoBERTw/o ${}_{\textrm{w/o G}}$ 9.31 2.99 2.50 2.65 2.96
fluffy OctoBERT 16.44 1.88 1.78 0.82 1.36 gym OctoBERT 5.13 2.14 0.44 0.74 0.69
OctoBERTw/o ${}_{\textrm{w/o G}}$ 15.61 1.83 1.71 1.37 1.47 OctoBERTw/o ${}_{\textrm{w/o G}}$ 4.88 3.73 1.30 1.08 1.45
circular OctoBERT 15.21 1.59 1.07 1.55 1.23 security OctoBERT 15.08 1.07 0.81 1.28 0.71
OctoBERTw/o ${}_{\textrm{w/o G}}$ 15.12 2.25 2.25 1.81 1.61 OctoBERTw/o ${}_{\textrm{w/o G}}$ 14.75 1.50 1.22 1.53 1.17
sink OctoBERT 14.23 1.17 0.92 1.11 1.38 cafe OctoBERT 6.28 1.90 1.38 1.98 1.39
OctoBERTw/o ${}_{\textrm{w/o G}}$ 15.49 1.84 1.65 1.60 1.84 OctoBERTw/o ${}_{\textrm{w/o G}}$ 7.03 2.17 1.92 2.08 1.72
doctor OctoBERT 13.03 1.17 1.05 1.38 1.18 teacher OctoBERT 16.68 1.95 2.15 1.52 1.48
OctoBERTw/o ${}_{\textrm{w/o G}}$ 12.44 1.17 1.23 1.39 1.58 OctoBERTw/o ${}_{\textrm{w/o G}}$ 16.08 2.68 2.37 1.85 1.83
foreign OctoBERT 9.48 0.62 0.95 0.85 0.47 student OctoBERT 16.28 1.38 1.07 1.20 1.03
OctoBERTw/o ${}_{\textrm{w/o G}}$ 10.01 1.03 0.88 1.18 0.95 OctoBERTw/o ${}_{\textrm{w/o G}}$ 16.52 2.21 1.29 1.40 1.61
diverse OctoBERT 16.44 0.60 0.22 0.52 0.24 newborn OctoBERT 16.43 1.71 0.88 0.91 1.11
OctoBERTw/o ${}_{\textrm{w/o G}}$ 16.05 0.81 0.65 0.97 0.65 OctoBERTw/o ${}_{\textrm{w/o G}}$ 16.30 2.02 1.32 1.61 1.76
product OctoBERT 10.25 0.84 0.75 1.39 1.15 pan OctoBERT 12.04 1.70 2.12 1.87 2.02
OctoBERTw/o ${}_{\textrm{w/o G}}$ 12.28 1.15 0.81 0.99 0.76 OctoBERTw/o ${}_{\textrm{w/o G}}$ 11.88 2.84 3.62 2.68 2.50
stove OctoBERT 16.15 2.63 2.64 1.94 2.72 telephone OctoBERT 14.09 1.18 0.96 1.05 0.96
OctoBERTw/o ${}_{\textrm{w/o G}}$ 16.13 3.06 4.30 3.08 2.98 OctoBERTw/o ${}_{\textrm{w/o G}}$ 13.42 1.17 1.50 1.46 1.38
steep OctoBERT 5.89 0.63 0.39 0.53 0.42 bamboo OctoBERT 14.54 2.02 1.20 0.76 1.02
OctoBERTw/o ${}_{\textrm{w/o G}}$ 7.30 1.46 2.42 0.87 1.93 OctoBERTw/o ${}_{\textrm{w/o G}}$ 15.40 3.01 1.38 1.09 1.42
warm OctoBERT 7.79 0.68 0.69 0.68 0.69 brush OctoBERT 11.17 1.88 2.13 1.81 2.45
OctoBERTw/o ${}_{\textrm{w/o G}}$ 8.67 1.05 1.01 0.79 0.85 OctoBERTw/o ${}_{\textrm{w/o G}}$ 13.69 2.51 2.89 2.39 2.83
aged OctoBERT 13.72 0.50 0.53 0.39 0.66 button OctoBERT 4.73 2.37 2.08 1.82 2.01
OctoBERTw/o ${}_{\textrm{w/o G}}$ 13.50 0.77 0.94 0.78 0.93 OctoBERTw/o ${}_{\textrm{w/o G}}$ 5.94 3.25 3.19 2.54 2.74
pizza OctoBERT 10.70 1.47 1.07 1.19 0.90
OctoBERTw/o ${}_{\textrm{w/o G}}$ 9.59 2.21 2.54 1.25 1.18